OpenStudy (ayeeeitskateee):

Please help, I need an explanation. Solve the equation. Check for extraneous solutions. 9|9-8x|=2x+3

OpenStudy (anonymous):

found Two solutions x=39/37 x=6/5

OpenStudy (anonymous):

9|-8x+9| = 2x+3 Clear the absolute-value bars by splitting the equation into its two cases, one for the Positive case and the other for the Negative case. The Absolute Value is 9|-8x+9| For the Negative case we'll use -9(-8x+9) For the Positive case we'll use 9(-8x+9) -9(-8x+9) = 2x+3 Multiply 72x-81 = 2x+3 Rearrange and Add up 70x = 84 Divide both sides by 70 x = (6/5) 9(-8x+9) = 2x+3 Multiply -72x+81 = 2x+3 Rearrange and Add up -74x = -78 Divide both sides by 74 -x = -(39/37) Multiply both sides by (-1) x = (39/37)

OpenStudy (anonymous):

^^

OpenStudy (ayeeeitskateee):

Yes, hold up. I'm reading it. Thanks!!

OpenStudy (ayeeeitskateee):

Okay, I think I got it! Thanks very much. I'm going to write this in my notebook.

OpenStudy (anonymous):

ok glad to help : )